OpenStudy (anonymous):

is there an easy way to find area in 6th grade math?

OpenStudy (rogue):

Area for what shape?

OpenStudy (anonymous):

baiscly square and rectangles

OpenStudy (anonymous):

some times triangles

OpenStudy (rogue):

For squares and rectangles, its always length times width. A triangle is really half of a rectangle if you rearrange it, so its area is 1/2 base * height.\[Rectangle: A = lw\]\[Square: A = lw = s^2\]\[Triangle: A = \frac {1}{2} bh\]

OpenStudy (anonymous):

that helps. thanks!

OpenStudy (rogue):

No problem, just do some practice with them and you'll easily learn them all =)

OpenStudy (anonymous):

hopefully

OpenStudy (unklerhaukus):

or get some unit cubes and see how many it takes to cover the surface of the shape